# Anchored Preference Optimization and Contrastive Revisions: Addressing Underspecification in Alignment

### v1

@@ Abstract
Large Language Models (LLMs) are often aligned using contrastive alignment objectives and preference pair datasets. The interaction between model, paired data, and objective makes alignment a complicated procedure, sometimes producing subpar results. We study this and find that (i) preference data gives a better learning signal when the underlying responses are contrastive, and (ii) alignment objectives lead to better performance when they specify more control over the model during training. Based on these insights, we introduce Contrastive Learning from AI Revisions (CLAIR), a data-creation method which leads to more contrastive preference pairs, and Anchored Preference Optimization (APO), a controllable and more stable alignment objective. We align Llama-3-8B-Instruct using various comparable datasets and alignment objectives and measure MixEval-Hard scores, which correlate highly with human judgments. The CLAIR preferences lead to the strongest performance out of all datasets, and APO consistently outperforms less controllable objectives. Our best model, trained on 32K CLAIR preferences with APO, improves Llama-3-8B-Instruct by 7.65%, closing the gap with GPT-4-turbo by 45%.
@@ 1 Introduction
In order to study the role of both (i) minimally contrastive preference data, and (ii) distinct alignment training dynamics, we individually align a model across four comparable preference datasets using five alignment objectives. One dataset is created through our CLAIR method. We compare this with two conventional judge-based datasets (Reinforcement Learning from AI Feedback; Bai et al. 2022). Finally, we consider an ablated version of CLAIR created to directly assess the impact of contrastiveness. We consider five distinct alignment objectives: DPO (Rafailov et al., 2024b), KTO (Ethayarajh et al., 2024), continued Supervised Fine-Tuning on the preferred answer, and two variants of our proposed APO. We measure MixEval-Hard accuracy (Ni et al., 2024) for each model, which correlates highly with model rankings produced by humans (Chiang et al., 2024).
@@ 2 Underspecification in Alignment
Given a collection of prompts $X$, a preference dataset is a set of triples $(x,y_{w},y_{l})$ , where $y_{w}$ and $y_{l}$ are, respectively, a winning (more preferred) and losing (less preferred) response to prompt $x$. The preference signal in such a dataset is essentially expressed by the difference between winning and losing outputs, illustrated in Figure 1 A. However, paired outputs can differ in many aspects, some of which are spurious and thus irrelevant to the preference. These spurious differences will generally create a challenging credit assignment problem. Outputs which are minimally contrastive differ along fewer axes, resulting in less spurious differences. Thus, the more contrastive preference pairs are, the more clear the learning signal becomes. Existing preference datasets vary meaningfully in their contrastiveness. For example, in the Stanford Human Preferences dataset (Ethayarajh et al., 2022), two outputs in a pair are simply responses to the same Reddit post, and thus they are not guaranteed to be especially comparable. An ideal preference dataset would consist of a very controlled difference between either example. This insight leads us to CLAIR (Section 3).
@@ 3 Contrastive Learning from Revisions
Part of the goal of Section 5 is to study the behavior of each of these datasets in the context of alignment efforts. However, one of the high-level goals of CLAIR is to generate examples that are minimally contrastive. We can assess this directly using some simple heuristics: the Jaccard similarity (token intersection over union) between $y_{w}$ and $y_{l}$ and the single-character Levenshtein edit distance between $y_{w}$ and $y_{l}$. The more contrastive a dataset is, the higher the Jaccard similarity should be and the lower the Levenshtein distance should be. Table 1 summarizes these analyses. By these measures, CLAIR delivers the most contrastive data by a wide margin.
@@ 5.1 Evaluation Methodology
Human judgments are ultimately the best indicator of how well a model is aligned with human preferences. Chatbot Arena (Chiang et al., 2024) uses thousands of pairwise human judgements to produce a ranking of model performance. However, collecting these judgments can be prohibitively expensive. To overcome this obstacle, we measure model performance through a benchmark which correlates highly with this Chatbot Arena ranking.
@@
We use the 2024-06-01 split of MixEval-Hard. Our evaluation of Llama-3-8B-Instruct before any additional alignment achieves a score of 41.45% on MixEval-Hard. The gap between Llama-3-8B-Instruct and GPT4-turbo is 17%.
Dataset Objective Max $\Delta$ Mean $\Delta$
Judge DPO 1.10 --0.74 (1.15)
off-policy KTO-pair --1.00 --2.89 (0.96)
SFT --1.95 --1.63 (1.06)
APO-zero 0.80 --1.99 (1.23)
APO-down 2.70 0.64 (0.98)
Judge DPO 4.00 0.56 (1.61)
on-policy KTO-pair 2.45 --0.51 (1.26)
SFT 0.65 --0.91 (1.01)
APO-zero 4.65 0.02 (1.66)
APO-down 3.65 1.60 (0.95)
CLAIR DPO 0.55 --1.68 (1.73)
KTO-pair 2.15 0.79 (0.98)
SFT 0.95 --1.63 (1.03)
APO-zero 7.65 2.93 (1.98)
APO-down --1.05 --5.22 (1.55)
Stronger DPO --5.00 --6.94 (1.03)
Preferred KTO-pair --1.20 --5.21 (1.27)
SFT 2.45 0.49 (1.31)
APO-zero --1.70 --2.72 (1.40)
APO-down --6.50 --12.51 (4.97)
Table 2: Max and mean MixEval-Hard improvements, aggregated over 18 epochs of aligning Llama-3-8B-Instruct on each dataset and alignment objective. Best overall performance bold, best performance per dataset underlined, standard deviation in parentheses. Our CLAIR dataset leads to the greatest overall performance improvement. Our controllable APO methods achieve the best performance across both Judged and CLAIR datasets. CLAIR with APO-zero closes the relative performance gap between Llama-3-8B-Instruct and GPT4-turbo by 45%.
@@ 5.3 Results
We report the maximal and mean MixEval-Hard improvement over all checkpoints from the same training run. This helps us understand both the best-case and average impact of alignment across the entire training procedure. This analysis is summarized in Table 2 for every dataset and objective; we now discuss these results in more detail.
@@ 5.3.1 Preference Data
To assess the quality of a particular dataset, we consider the performance of that dataset when paired with its best objective. Using the APO-zero objective, the contrastive CLAIR dataset leads to the greatest improvement. Specifically, CLAIR leads to the greatest maximal improvement of +7.65% and the greatest average improvement of +2.93% out of all our experiments. This improvement of +7.65% closes the relative gap with GPT4-turbo by 45% using only 32K pairs.
@@
A breakdown of performance in function of MixEval-Hard’s constituent benchmarks is given in Appendix B for every dataset.
@@ 5.3.2 Alignment Objectives
Anchored Preference Optimization (APO) consistently leads to the greatest performance increase for every preference dataset, with the exception of the Stronger Preferred dataset, where all contrastive objectives deteriorate the model. The relation between the preference dataset and the target model controls which variant of APO is best for any dataset, as predicted in Section 2. APO-down results in the best performance when winning outputs are generally worse than the target model, as is the case for the off-policy judge dataset. APO-zero is the best objective when winning outputs are generally better than the target model, as is the case for CLAIR and on-policy judge datasets. The difference between alignment objectives is less salient for the on-policy judge dataset as compared to CLAIR, since winning on-policy judge outputs are only slightly better than Llama-3-8B-Instruct on average. Winning CLAIR outputs may be vastly better than Llama-3-8B-Instruct since they are produced by a stronger model, making the different in alignment objectives more noticeable.
@@ Acknowledgements
We thank Kawin Ethayarajh, Eugen Hotaj, and Nathan Lambert for their feedback. We thank Stas Bekman for his help and support.

### v2

@@ Abstract
Large Language Models (LLMs) are often aligned using contrastive alignment objectives and preference pair datasets. The interaction between model, paired data, and objective makes alignment a complicated procedure, sometimes producing subpar results. We study this and find that (i) preference data gives a better learning signal when the underlying responses are contrastive, and (ii) alignment objectives lead to better performance when they specify more control over the model during training. Based on these insights, we introduce Contrastive Learning from AI Revisions (CLAIR), a data-creation method which leads to more contrastive preference pairs, and Anchored Preference Optimization (APO), a controllable and more stable alignment objective. We align Llama-3-8B-Instruct using various comparable datasets and alignment objectives and measure MixEval-Hard scores, which correlate highly with human judgments. The CLAIR preferences lead to the strongest performance out of all datasets, and APO consistently outperforms less controllable objectives. Our best model, trained on 32K CLAIR preferences with APO, improves Llama-3-8B-Instruct by 7.65%, closing the gap with GPT4-turbo by 45%.
@@ 1 Introduction
In order to study the role of both (i) minimally contrastive preference data, and (ii) distinct alignment training dynamics, we individually align a model across four comparable preference datasets using five alignment objectives. One dataset is created through our CLAIR method. We compare this with two conventional judge-based datasets (Reinforcement Learning from AI Feedback; Bai et al. 2022). Finally, we consider an ablated version of CLAIR created to directly assess the impact of contrastiveness. We consider five distinct alignment objectives: DPO (Rafailov et al., 2024b), KTO (Ethayarajh et al., 2024), continued Supervised Fine-Tuning on the preferred answer, and two variants of our proposed APO. We measure MixEval-Hard accuracy (Ni et al., 2024) and length-controlled AlpacaEval scores (Dubois et al., 2024) for each model, both correlate highly with model rankings produced by humans (Chiang et al., 2024).
@@ 2 Underspecification in Alignment
Given a collection of prompts $X$, a preference dataset is a set of triples $(x,y_{w},y_{l})$ , where $y_{w}$ and $y_{l}$ are, respectively, a winning (more preferred) and losing (less preferred) response to prompt $x$. The preference signal in such a dataset is essentially expressed by the difference between winning and losing outputs, illustrated in Figure 1 A. However, paired outputs can differ in many aspects, some of which are spurious and thus irrelevant to the preference. These spurious differences will generally create a challenging credit assignment problem. Outputs which are minimally contrastive differ along fewer axes, resulting in less spurious differences. Thus, if preference pairs produce a clearer minimal contrast, the alignment learning signal becomes more clear. Existing preference datasets vary meaningfully in their contrastiveness. For example, in the Stanford Human Preferences dataset (Ethayarajh et al., 2022), two outputs in a pair are simply responses to the same Reddit post, and thus they are not guaranteed to be especially comparable. An ideal preference dataset would consist of a very controlled difference between either example. This insight leads us to CLAIR (Section 3).
@@ 3 Contrastive Learning from Revisions
Part of the goal of Section 5 is to study the behavior of each of these datasets in the context of alignment efforts. However, one of the high-level goals of CLAIR is to generate examples that are minimally contrastive. We can assess this directly using some simple heuristics: the Jaccard similarity (token intersection over union) between $y_{w}$ and $y_{l}$ and the single-character Levenshtein edit distance between $y_{w}$ and $y_{l}$. The dataset with better minimal contrasts should result in a higher Jaccard similarity and a lower Levenshtein distance. Table 1 summarizes these analyses. By these measures, CLAIR delivers the best contrastive data by a wide margin.
@@ 5.1 Evaluation Methodology
Human judgments are ultimately the best indicator of how well a model is aligned with human preferences. Chatbot Arena (Chiang et al., 2024) uses thousands of pairwise human judgements to produce a ranking of model performance. However, collecting these judgments can be prohibitively expensive. To overcome this obstacle, we measure model performance through benchmarks which correlate highly with this Chatbot Arena ranking.
@@
Our evaluation of Llama-3-8B-Instruct before any additional alignment achieves a score of 41.45% on the 2024-06-01 version of MixEval-Hard. The gap between Llama-3-8B-Instruct and GPT4-turbo is 17%. On the 2024-08-11 split, Llama-3-8B-Instruct achieves 40.5%.
Additionally, we consider the length-controlled LC-AlpacaEval2.0 win rate (Dubois et al., 2024) as an auxiliary evaluation method. LC-AlpacaEval2.0 measures how often a generated output is preferred to a reference output, as judged by a GPT4 model across a wide set of instructions. On this benchmark, Llama-3-8B-Instruct attains a win rate of 22.92% whereas GPT4 has a win rate of 50%.
ME-Hard 2024-06-01 ME-Hard 2024-08-11 LC-AlpacaEval2.0
Dataset Objective Max $\Delta$ Mean $\Delta$ Max $\Delta$ Mean $\Delta$ Score $\Delta$ Length $\Delta$
Judge DPO 1.10 --0.74 (1.15) 4.30 2.85 (0.75) 2.94 --158
off-policy KTO-pair --1.00 --2.89 (0.96) 4.05 1.18 (1.67) --5.69 --437
SFT --1.95 --1.63 (1.06) 2.85 0.42 (1.20) --22.29 12669
APO-zero 0.80 --1.99 (1.23) 4.65 1.26 (1.62) --2.42 --395
APO-down 2.70 0.64 (0.98) 4.80 3.52 (0.85) 2.40 --203
Judge DPO 4.00 0.56 (1.61) 5.20 2.71 (1.41) 4.98 341
on-policy KTO-pair 2.45 --0.51 (1.26) 5.05 1.13 (1.70) 3.02 452
SFT 0.65 --0.91 (1.01) 4.20 2.55 (0.70) 1.34 156
APO-zero 4.65 0.02 (1.66) 5.35 2.19 (1.28) 5.51 484
APO-down 3.65 1.60 (0.95) 4.25 3.06 (0.76) 7.63 386
CLAIR DPO 0.55 --1.68 (1.73) 5.05 2.77 (1.40) 2.65 966
KTO-pair 2.15 0.79 (0.98) 4.65 2.92 (0.86) 4.33 160
SFT 0.65 --0.91 (1.01) 2.70 0.92 (1.21) --0.47 6108
APO-zero 7.65 2.93 (1.98) 5.95 4.39 (0.89) 5.08 520
APO-down --1.05 --5.22 (1.55) --1.20 --3.61 (1.05) --6.30 2559
Stronger DPO --5.00 --6.94 (1.03) --3.10 --4.40 (0.98) --2.89 597
Preferred KTO-pair --1.20 --5.21 (1.27) 2.25 0.50 (1.13) 0.71 153
SFT 2.45 0.49 (1.31) 5.05 2.73 (1.21) 6.99 1883
APO-zero --1.70 --2.72 (1.40) --4.85 --12.02 (5.38) 0.89 243
APO-down --6.50 --12.51 (4.97) 1.65 0.16 (1.22) 1.87 10001
Table 2: Max and mean MixEval-Hard improvements for 2024-06-01 and 2024-08-11 splits, aggregated over 18 epochs of aligning Llama-3-8B-Instruct. Best overall performance bold, best performance per dataset underlined, standard deviation in parentheses. While MixEval-Hard correlates better with human model ratings, we also report the average LC-AlpacaEval2.0 score increase over the two best MixEval-Hard checkpoints, and average length increase (in characters) of the responses. CLAIR leads to the greatest overall performance improvement on MixEval-Hard. APO methods achieve the best performance across both Judged and CLAIR datasets.
@@ 5.3 Results
We report the maximal and mean MixEval-Hard improvement over all checkpoints from the same training run. This helps us understand both the best-case and average impact of alignment across the entire training procedure. We use both 2024-06-01 and 2024-08-11 versions of MixEval-Hard, which each feature a distinct set of queries. Due to the increased cost associated with LC-AlpacaEval2.0, we only measure the win rate for the two best MixEval-Hard checkpoints and report their average. This analysis is summarized in Table 2 for every dataset and objective; we now discuss these results in more detail.
@@ 5.3.1 Preference Data
To assess the quality of a particular dataset, we consider the performance of that dataset when paired with its best objective. Using the APO-zero objective, the contrastive CLAIR dataset leads to the greatest improvement. On the 2024-06-01 split of MixEval-Hard, CLAIR leads to the greatest maximal improvement of +7.65% and the greatest average improvement of +2.93% out of all our experiments. This improvement of +7.65% closes the relative gap with GPT4-turbo by 45% using only 32K pairs.
@@
For the LC-AlpacaEval2.0 benchmark, on-policy judge leads to better performance than CLAIR. However, LC-AlpacaEval2.0 correlates less with human-produced model rankings than MixEval-Hard.
A breakdown of performance in function of MixEval-Hard’s constituent benchmarks is given in Appendix B for every dataset. We remark on a potential issue we have encountered with LC-AlpacaEval2.0 in Appendix C.
@@ 5.3.2 Alignment Objectives
On MixEval-Hard, Anchored Preference Optimization (APO) consistently leads to the greatest performance increase for every preference dataset, with the exception of the Stronger Preferred dataset, where all contrastive objectives deteriorate the model. The relation between the preference dataset and the target model controls which variant of APO is best for any dataset, as predicted in Section 2. APO-down results in the best performance when winning outputs are generally worse than the target model, as is the case for the off-policy judge dataset. APO-zero is the best objective when winning outputs are generally better than the target model, as is the case for CLAIR and on-policy judge datasets. The difference between alignment objectives is less salient for the on-policy judge dataset as compared to CLAIR, since winning on-policy judge outputs are only slightly better than Llama-3-8B-Instruct on average. Winning CLAIR outputs may be vastly better than Llama-3-8B-Instruct since they are produced by a stronger model, making the different in alignment objectives more noticeable.
@@ Acknowledgements
We thank Kawin Ethayarajh, Eugen Hotaj, and Nathan Lambert for their feedback. We thank Stas Bekman for his help and support. KD gratefully acknowledges funding from the FWO Fundamental Research PhD Fellowship (11632223N).
@@ Appendix B MixEval-Hard Performance Breakdown
Appendix C How well does AlpacaEval control for response lengths?
GPT4 is known to favor more verbose responses, which can artificially inflate AlpacaEval win rates for verbose models (Dubois et al., 2024). To counteract this bias, Dubois et al. (2024) estimate a length-controlled AlpacaEval win rate, which we report on in Table 2. Specifically, the authors adopt a causal inference framework to answer the question "What would the AlpacaEval metric be, if the outputs of all models had the same length as those of the baseline?" (Dubois et al., 2024).
In order to meaningfully apply causal inference, a few key assumptions need to be met. The Positivity assumption (Hernán and Robins, 2006) states that, when estimating the effect of a treatment (e.g., the effect of response length on a model’s win rate), for all covariates (e.g., for all models), there are at least some subjects which receive the treatment (e.g., there are some responses of that model which are long and some which are not). Intuitively, the Positivity assumption applied to the length-control question states that you need to observe at least some long and some short responses for every model in order to accurately estimate how the response length influences the model’s win rate.
The AlpacaEval framework does not check if this Positivity assumption is met, potentially giving bad estimates for the length-controlled win rates in some settings. If a certain model consistently generates responses longer than those of the baseline, it is impossible to accurately estimate how good the responses would be if they were as long as the baseline.
This may give us insights into some of our length-controlled AlpacaEval win rates. For example, the SFT result on the Stronger Preferred dataset in Table 2 seems disproportionately high in comparison to the MixEval-Hard results for that same experiment. This model is considerably more verbose than Llama-3-8B-Instruct, as evident from the large response length increase associated with this experiment ( + 1883 characters on average). It is possible the Positivity constraint was not met for this experiment, causing the length-controlled framework of AlpacaEval to provide inaccurate estimates.
One potential avenue towards a more robust length-controlled win rate would be to specifically prompt models to generate shorter or longer answers if the Positivity constraint is not met. Unfortunately, a more thorough study of length-controlled win rate is out of scope for this work.